A balloon Rises with uniform velocity of 10 metre per second when the balloon is at a height of 72.5metre if a stone is dropped from the time taken to reach the ground is

A  balloon Rises with uniform velocity of 10 metre per second when the balloon is at a height of 72.5metre if a stone is dropped from the time taken to reach the ground is

Grade:9

1 Answers

Pooja
4 years ago
Here the balloon is rising with uniform velocity
v=10m/s
At height h=72.5m , a stone is dropped
The time consumed till now can be found by
h=v.t
72.5=10×t
So t =7.25 sec
 
Now the ball will have initial velocity suppose V
Therefore V=10m/s
Hence it will ascend till it's velocity becomes zero and then fall on ground
 
Time for ball to ascend(t')
V=gt'
10=10×t'
Therefore t'=1sec
 
Now the extra height H acheived by ball
V^2=2gh
10×10 =2×10 × H
Therefore H =5m
Now the ball is at highest point of height (h')72.5+5=77.5m
Time to reach ground from highest point (t'')can be calculated by
t''=sqrt (2h'/g)
   =sqrt(2×77.5/10)
   =4 sec(approx.)
 
So total time taken by ball to reach ground from the point of dropping is t' +t''
Time taken=1 + 4 =5sec
 
And time taken by ball from the starting of balkon rising is 5+7.25=7.75 sec